if the line passing through the points (a,1) and (1,6) is parallel to the line passing through the points (-7,8) and (a+2,1) what is the value of a?
Answer by Boreal(15235)

Parallel lines have the same slope
the slope is difference in y divided by difference in x. For the first pair, it is
5/(1-a) = (-7)/(a+2+7), which is (a+9)
Cross-multiply
(-7)(1-a)=5(a+9)
-7+7a=5a+45
2a=52
a=26 ANSWER
(26, 1) and (1, 6) which has a slope of 5/-25 or -1/5
(-7, 8) and (28, 1) which has a slope of -7/25 or -1/5